Which of the following sets have the same cardinality? Select all that apply.


LaTeX: \mathbb{N} N

[0,1]

LaTeX: \mathbb{R} R

(0,1)
1
Expert's answer
2020-07-13T19:17:52-0400

Cardinality of "\mathbb{R}" is the same as cardinality of (0,1), because there is a bijection from first set to second: "\dfrac{1}{1+e^{-x}}"


Cardinality of [0,1] is the same as cardinality of (0,1), because there is a bijection from first set to second:

From the first set cut a sequence of points "0,1/4, 1/4^2, 1/4^3, \dots" and insert them to coordinates "1/4, 1/4^2, 1/4^3, 1/4^4, \dots" respectively. It is a bijection from [0,1] to (0,1].

Then set cut a sequence of points "1,(1-1/4), (1-1/4^2), (1-1/4^3), \dots" and insert them to coordinates "(1-1/4), (1-1/4^2), (1-1/4^3), (1-1/4^4) \dots" respectively. It is a bijection from (0,1] to (0,1).


Cardinality of "\mathbb{N}" is less then cardinality of "\mathbb{R}", because "\mathbb{N}" is a countable set and "\mathbb{R}" is a continuum. "\mathbb{R}" has the same cardinality as "2^\mathbb{N}".